An electron is accelerated through a potential difference off 100 volts, it's velocity is nearly:
A. 1.9 x 10m/sec
B. 1.9 x 10m/sec
C. 3.8 x 10m/sec
D. 3.8 x 10m/sec
 

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
Potential difference of the electron = 1000 volts (Given)
Potential difference is the difference of an electrical potential between two points. If the voltage is given as v, then the energy of electron will be 
= 1/2mv² 
= ev = v = √2eV/m
Thus, 
= √ 2 × 1.6 × 10`19 × 1000 / 9.1 × 10`31
= 1.875 × 10⁷
= 1.9 × 10⁷ m/s
Therefore, if an electron is accelerated through a potential difference of 1000 volts. Its velocity is nearly 1.9 × 10⁷ m / s.